# DESIGNER: Design-Logic-Guided Multidisciplinary Data Synthesis for LLM Reasoning

### v5

@@ Appendix D Baseline Datasets
Table 5: The baseline datasets utilized in our experiments. For the Nemotron-Post-Training-v1 dataset, only the data from the math and STEM categories was used.
@@
OpenThoughts3
https://huggingface.co/datasets/open-thoughts/OpenThoughts3-1.2M
Nemotron-Post-Training-v1
https://huggingface.co/datasets/nvidia/Nemotron-Post-Training-Dataset-v1
WebInstruct (Full)
https://huggingface.co/datasets/TIGER-Lab/WebInstructFull
NaturalReasoning
https://huggingface.co/datasets/facebook/natural_reasoning
@@ Appendix H Experimental Details
MMLU (hendrycks2020measuring) 57 disciplines including STEM, humanities, social sciences, and professional fields 1
MMLU-Pro (wang2024mmlu) Math, Physics, Chemistry, Law, Engineering, Economics, Health, Psychology, Business, Biology, Philosophy, Computer Science, History, Other 1
GPQA-Diamond (rein2024gpqa) Physics, Chemistry, Biology 10
GPQA-Main (rein2024gpqa) Physics, Chemistry, Biology 10
SuperGPQA (du2025supergpqa) 285 graduate-level disciplines across 13 fields (e.g., Engineering, Management, Economics, Education, History, Science, Medicine, Law, Sociology, Philosophy, Agriculture, Literature) 1
@@ Appendix I Performance on GSM8K and MATH-500
Notably, models trained on either DLR-Book or DLR-Web+Book achieve 95.60% on GSM8K, matching the official Qwen3-8B score. GSM8K performance is already near saturation. Prior work (vendrow2025large) reports that approximately 5% of GSM8K questions are mislabeled or ambiguous, which makes 100% accuracy unattainable and positions 95.60% near the practical upper bound. Even stronger LLMs fall within a similar range. For instance, according to the Llama 3 technical report (dubey2024llama), GPT-4o reaches 96.1 on GSM8K under 8-shot CoT setting.

### v6

@@ Appendix D Baseline Datasets
Table 5: The baseline datasets utilized in our experiments. For the Nemotron-Post-Training-v1 dataset, we used the data from the math and STEM categories.
@@
OpenThoughts3 (Guha et al., 2025) https://huggingface.co/datasets/open-thoughts/OpenThoughts3-1.2M
Nemotron-Post-Training-v1 (Bercovich et al., 2025) https://huggingface.co/datasets/nvidia/Nemotron-Post-Training-Dataset-v1
WebInstruct (Full) (Yue et al., 2024) https://huggingface.co/datasets/TIGER-Lab/WebInstructFull
NaturalReasoning (Yuan et al., 2025) https://huggingface.co/datasets/facebook/natural_reasoning
@@ Appendix H Experimental Details
MMLU (Hendrycks et al., 2020) 57 disciplines including STEM, humanities, social sciences, and professional fields 1
MMLU-Pro (Wang et al., 2024b) Math, Physics, Chemistry, Law, Engineering, Economics, Health, Psychology, Business, Biology, Philosophy, Computer Science, History, Other 1
GPQA-Diamond (Rein et al., 2024) Physics, Chemistry, Biology 10
GPQA-Main (Rein et al., 2024) Physics, Chemistry, Biology 10
SuperGPQA (Du et al., 2025) 285 graduate-level disciplines across 13 fields (e.g., Engineering, Management, Economics, Education, History, Science, Medicine, Law, Sociology, Philosophy, Agriculture, Literature) 1
@@ Appendix I Performance on GSM8K and MATH-500
Notably, models trained on either DLR-Book or DLR-Web+Book achieve 95.60% on GSM8K, matching the official Qwen3-8B score. GSM8K performance is already near saturation. Prior work (Vendrow et al., 2025) reports that approximately 5% of GSM8K questions are mislabeled or ambiguous, which makes 100% accuracy unattainable and positions 95.60% near the practical upper bound. Even stronger LLMs fall within a similar range. For instance, according to the Llama 3 technical report (Dubey et al., 2024), GPT-4o reaches 96.1 on GSM8K under 8-shot CoT setting.